OpenStudy (anonymous):

Easy Algebra Question: A 9th grade math class has 27 students. There are twice as many girls in this class as there are boys. Write an system of equations that could be used to model this situation. (WILL FAN AND MEDAL ;))

OpenStudy (hysusonic):

so do you know what you can do?

OpenStudy (anonymous):

Not really

OpenStudy (hysusonic):

well first throw out the useless information and keep the useful things such as the total amount of students and how many boys there are to girls

OpenStudy (hysusonic):

27 students twice as many girls than boys

OpenStudy (hysusonic):

you with me so far?

OpenStudy (anonymous):

Yes

OpenStudy (hysusonic):

okay so you know you your answer has to add up to 27 and you have 2 variables, what can you make with this?

OpenStudy (anonymous):

divide something ?

OpenStudy (hysusonic):

nope add the 2 variable which would make x+y=27

OpenStudy (anonymous):

Ok

OpenStudy (hysusonic):

so we have our first system

OpenStudy (anonymous):

Right

OpenStudy (hysusonic):

do you know how to get the second one?

OpenStudy (anonymous):

No

OpenStudy (hysusonic):

well we would have to find the y-value (the number of girls)

OpenStudy (anonymous):

Ugh I have no idea how to do this

OpenStudy (hysusonic):

so you set an equation to Y which would be x+x=y or 2x=y

OpenStudy (anonymous):

ok

OpenStudy (hysusonic):

wait no i'm confused now... whoops sorry this is where i have to leave you but maybe @misty1212 or @Awolflover1 can help

OpenStudy (mathstudent55):

The first equation is x + y = 27 For the second equation, there are twice as many girls (y) as boys (x), so you write: y = 2x Now you have a system of equations: x + y = 27 y = 2x